OpenStudy (fanduekisses):

How to find the exact value using the Half-Angle Identity: 7pi/8 I am stuck

OpenStudy (solomonzelman):

For what trigonometric function? sine cosine or what ?

OpenStudy (fanduekisses):

tan

OpenStudy (fanduekisses):

So far I get = \[\frac{ 1-\sqrt{2} }{ 1+\sqrt{2} }\]

OpenStudy (fanduekisses):

is that right?

OpenStudy (solomonzelman):

tan(157.5 )= tan (315 / 2 ) \(\Huge\color{blue}{ \bf ± \sqrt{ \frac{1-cos(a)}{1+cos(a)} } }\) I need to check...

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf ± \sqrt{ \frac{1-cos(315)}{1+cos(315)} } }\) cos (315)= cos(360-45)=cos 360 cos 45 - sin 360 sin 45 = cos 360=1 sin 360 = 0 cos 360 cos 45 - sin 360 sin 45 = 1 * cos 45 - 0 * sin 45 = cos 45 cos45=1/2  \(\Huge\color{blue}{ \bf ± \sqrt{ \frac{1-1/2}{1+1/2} } }\)

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf ± \sqrt{ \frac{1-1/2}{1+1/2} } }\) \(\Huge\color{blue}{ \bf ± \sqrt{ \frac{2/2-1/2}{2/2+1/2} } }\) \(\Huge\color{blue}{ \bf ± \sqrt{ \frac{1/2}{3/2} } }\) \(\Huge\color{blue}{ \bf ± \sqrt{ \frac{1}{2} \div \frac{3}{2} } }\) \(\Huge\color{blue}{ \bf ± \sqrt{ \frac{1}{2} \times \frac{2}{3} } }\) \(\Huge\color{blue}{ \bf ± \sqrt{ \frac{1}{3} } }\)

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf ± \sqrt{ \frac{1}{3} } }\) \(\Huge\color{blue}{ \bf ± \frac{ \sqrt{1}}{ \sqrt{3}} }\) \(\Huge\color{blue}{ \bf ± \frac{ 1}{ \sqrt{3}} }\) \(\Huge\color{blue}{ \bf ± \frac{ \sqrt{3}}{ 3} }\) is the ans.

OpenStudy (fanduekisses):

but it says the answers can only be 1 + square root of two 1 - square root of two -1 + square root of two -1 - square root of two